# Towards Proof Repair in Cubical Agda

### 02 Feb 2024 v2

@@ Abstract.
Recent work introduced an algorithm and tool in Coq to automatically repair broken proofs in response to changes that correspond to type equivalences. We report on case studies for manual proof repair across type equivalences using an adaptation of this algorithm in Cubical Agda. Crucially, these case studies capture proof repair use cases that were challenging to impossible in prior work in Coq due to type theoretic limitations, highlighting three benefits to working in Cubical Agda: (1) quotient types enrich the space of repairs we can express as type equivalences, (2) dependent path equality makes it possible to internally state and prove correctness of repaired proofs relative to the original proofs, and (3) functional extensionality and transport make it simple to move between slow and fast computations after repair. They also highlight two challenges of working in Cubical Agda, namely those introduced by: (1) lack of tools for automation, and (2) proof relevance, especially as it interacts with definitional equality. We detail these benefits and challenges in hopes to set the stage for later work in proof repair bridging the benefits of both languages.
Proof Repair, Cubical Agda
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2024, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Proof repair (Ringer, 2021) aims to simplify this process by introducing algorithms and tools that fix formal proofs in response to breaking changes. In this paper, we report on manual proof repair case studies in a fragment of Cubical Agda. Our focus is on a particular class of changes: those that can be described by type equivalences, as in the Pumpkin Pi Coq plugin (Ringer et al., 2021). Using the features of Cubical Agda, however, we are able to construct equivalences between types that cannot naturally be expressed in Coq. We investigate how the proof repair algorithm from Pumpkin Pi carries over to Cubical Agda on two case studies:
@@
    We examine two non-isomorphic representations of the integers: (a) the representation used in the Cubical Agda standard library, and (b) the Grothendieck group completion of the natural numbers. Cubical Agda allows us to formally construct the latter, which would be challenging in Coq, and repair the addition function and several proofs across representations.
@@
    We implement two variations of queues: (a) those backed by lists and (b) those backed by pairs of lists. These types are not equivalent, but by leveraging Cubical Agda’s quotient types, we can construct an equivalence that describes the change and repair functions and proofs across it.
Cubical Agda (Vezzosi et al., 2019) extends Agda to give computational meaning to the univalence axiom and higher inductive types by way of cubical type theory (Angiuli et al., 2017; Cohen et al., 2018), enabling for a computational notion of functional extensionality and quotient types. In moving to Cubical Agda, we report on what is gained:
@@
    Quotient Equivalences: We can support equivalences between quotient types, thereby supporting a broader class of changes than those handled by prior work. Recent work in Cubical Agda showed that certain relations describing changes in behavior can be adjusted to equivalences between quotient types (Angiuli et al., 2021). Since Coq lacks quotient types, prior work in Coq could support only a more restricted class of changes: those that can be described by equivalences between sigma types. By taking advantage of cubical’s quotient types, we broaden the class of changes we can support.
@@
    Internal Correctness: We can state and prove the correctness of repaired proofs internally in Cubical Agda, using cubical’s dependent path equality. Since Coq lacks univalence, prior work in Coq was resigned to stating the correctness theorems on paper inside of a univalent metatheory. By taking advantage of cubical’s rich notion of path equality, we are able to state and prove previously unproven theorems directly and formally inside of Cubical Agda.
@@
    Repair to Fast Functions: We can move between fast and slow representations in Cubical Agda much more easily than in Coq. When repairing proofs across equivalences, the repair methodology we adopt from previous work imposes the same inductive structure on the repaired functions and proofs as the original. This in combination with Coq’s type theory is why prior work used ad hoc methods to move between, say, inefficient (but easy to reason about) unary addition, and efficient (but difficult to reason about) binary addition, even when one had a principled way to repair from unary to binary representations of the naturals. In Cubical Agda, however, functional extensionality is a theorem, so we can take proof repair one step further, replacing slow repaired functions with fast ones by chaining proof repair with transport across the extensional equality.
We also report on the challenges we experienced:
@@
    Automation: Automation in Cubical Agda is difficult due to both engineering limitations and type theoretic challenges. Because of this, our proof repair case studies in Cubical Agda are still manual. In contrast, Coq’s rich plugin system made it possible to build powerful external proof repair automation in prior work.
@@
    Proof Relevance: While Cubical Agda makes it possible to prove internal correctness of certain repaired proofs for the first time, doing so can be challenging thanks to Cubical Agda’s proof relevance—the fact that it matters not just what is proven, but also how that is done. Specifically, we struggled to prove that concrete proofs were repaired correctly, even when we were able to show that all of the components from which they were composed were repaired correctly. This is because, when it came to composing internal correctness proofs, the contents of those individual components had to line up. Aligning proofs about repaired terms with proofs about repaired types in a way that had the right compositional and definitional behavior proved at times to be prohibitively difficult.
All of our code is available. [^1] https://github.com/InnovativeInventor/proof-repair-cubical Our goal is to set the stage for proof repair work bridging the two different type theories and tools, in hopes to get the best of both worlds.
@@ 2. Problem Definition
We move this problem into a fragment of cubical type theory. By working in cubical, we can expand the set of pairs of equivalent types $A$ and $B$ that we can repair proofs over to include an extremely useful set of changes in types not describable in Coq—quotient type equivalences (Section 2.1). Furthermore, we can internally and formally state what it means for any given repaired proof defined over $B$ to correctly correspond to the original proof over $A$—in terms of dependent path equality (Section 2.2). We will ground this in concrete examples with our case studies in Section 4.
@@ 2.1. Scope: Quotient Type Equivalences
We wish to generalize Pumpkin Pi to Cubical Agda in order to capture broader use cases. In this paper, we perform proof repair across quotient type equivalences. We will describe what makes these quotient type equivalences soon. But in general, a type equivalence between two types $A$ and $B$ is an isomorphism (a pair of functions that are mutual inverses) satisfying a particular coherence property (Univalent Foundations Program, 2013). It is possible to form an equivalence from any isomorphism, for example by using isoToEquiv in Cubical Agda:
  isoToEquiv : Iso A B normal-$\to$ A $\simeq$ B
where $\simeq$ in Cubical Agda denotes a type equivalence.
@@
Like prior work (Ringer et al., 2021), we consider changes in datatypes that can be described by these type equivalences. But by working in Cubical, we can support a broader class of changes than type equivalences in Coq can encode. The trick is to encode changes in datatypes as equivalences between quotient types—types equipped with an equivalence relation describing what makes two elements of that type “the same.”
Quotient types are supported natively in cubical type theory. In Cubical Agda, quotient types (denoted $A/R$ for a type $A$ and relation $R$) are encoded as higher inductive types taking a type and a relation:
An element of this quotient is the equivalence class of any $a$ of type $A$ (denoted $[a]$):
    [_] : (a : A) normal-$\rightarrow$ A / R
Given two elements $[a_{1}]$ and $[a_{2}]$, we have that $[a_{1}]\equiv[a_{2}]$ whenever $R\ a_{1}\ a_{2}$:
[…]
even when it is not true that $a_{1}\equiv a_{2}$.
To give a familiar example of a quotient type, we can construct the natural numbers mod 2, $\mathbb{N}/2$. Two naturals are in the same equivalence class if they have the same parity. We can capture this with the following type:
    ~ : $\mathbb{N}$ $\to \mathbb{N}$ normal-$\to$ Type
    ~ n m = (n mod 2) $\equiv$ (m mod 2)
    $\mathbb{N}/2$ = $\mathbb{N}$ / $\sim$
@@
The resulting type has two equivalence classes, and so every element of the type is equal to either [ 0 ] or [ 1 ]. For another element, like [ 2 ], we have a proof of equality between [ 0 ] and [ 2 ] via eq/ [ 0 ] [ 2 ] refl. This resulting type is then isomorphic to the booleans via the map [ 0 ] $\mapsto$ false, [ 1 ] $\mapsto$ true, and this can be used to construct a type equivalence via isoToEquiv. In Section 4, when we discuss our case studies, we will discuss more complicated examples of quotient types and equivalences.
In this paper, we further restrict ourselves to equivalences between set quotients, which have one additional constraint:
@@
That is, all proofs of equality between elements of the quotient must themselves be equal to each other. In homotopy type theory parlance, the type $A/R$ must be an h-set; in broader dependent type theory parlance, uniqueness of identity proofs must provably hold for the type $A/R$ [^2] While homotopy type theory makes it possible to define types for which identity proofs are not unique, reasoning about equalities between those types can be challenging, so we defer reasoning about them to later work.
Quotient types are commonly used for mathematics—a use case we will highlight in Section 4.1. But they also let us cleanly express changes in the implementation of datatypes—a benefit we will highlight in Sections 4.2 and 5.1.1.
Repairing proofs across changes that can be expressed as quotient type equivalences amounts to instantiating the proof term transformation from Pumpkin Pi inside of a fragment of cubical type theory; we describe this in Section 3.
@@ 2.2. Goal: Dependent Path Equality
In cubical type theory, we can state what it means for a repaired proof to be correctly related to the original proof. This is due to univalence (Cohen et al., 2018; Angiuli et al., 2017) in cubical type theory—equivalence is equivalent to equality (Univalent Foundations Program, 2013). [^3] In cubical type theory, univalence is not an axiom, but rather follows computationally from more primitive constructs. We take advantage not of univalence directly, but of these more primitive constructive constructs. In contrast, prior proof repair work in Coq could describe correctness externally in some univalent metatheory (Ringer et al., 2021), but could not generally prove or state correctness internally inside of Coq without the axiom of functional extensionality (Tabareau et al., 2021). This highlights another benefit of working in Cubical, detailed in Section 5.
We define correctness in terms of a generalization of path equality, the primitive notion of propositional equality in cubical type theory. In Cubical Agda, if two terms a and b of the same type A are path equal to each other, we can produce a term of type:
  a $\equiv$ b
Our two-element types from before provide an example of path equality. Using univalence, the isomorphism between $\mathbb{N}/2$ and the booleans can be turned into an equality:
  $\mathbb{N}/2$equiv bool
Path equality is actually an instantiation of the more powerful notion of dependent path equality, for equality between terms with different types (that is, heterogeneous equality). In Cubical Agda, two terms a of type A and b of type B are dependently path equal if we can produce a term of type:
@@
where p is a path between A and B. Non-dependent path equality $\equiv$ is defined in terms of dependent path equality at the identity path from a type to itself. For example, let p be our proof that $\mathbb{N}/2$equiv bool. Then the type:
@@
We can state correctness of a repaired proof relative to the corresponding original version by stating the correct PathP type—the one between the original term and the repaired term at the path between their types. Furthermore, we can do this directly inside of Cubical Agda, without any axioms. The type PathP p [ 0 ] false we described above gives such an example: if we repair [ 0 ] to false across $\mathbb{N}/2\equiv$ bool, then this type states that this repair was correct. We describe how to define correctness more generally using PathP when we detail our approach in Section 3.
There are two important things to note here. The first is that, while such a PathP type can in theory be not just stated in Cubical Agda but also proven (inhabited), in practice, finding a term with the correct type can be quite difficult (see Section 5). The second is that correctness of a repaired proof relies not only on the right PathP type being inhabited, but also on the new version of the proof no longer containing references to the old version of the proof. This we do not state internally to Cubical Agda, but it should still hold metatheoretically for repair to be correct (Ringer et al., 2021; Ringer, 2021). This is also the primary way in which proof repair in cubical type theory diverges from (dependent) transport: using transport to move proofs across equalities forces both equivalent types to eternally remain in the codebase.
3. Approach
We detail our approach to performing proof repair in Cubical Agda. To begin, we recall the algorithm for proof repair across equivalences in Coq that we will adapt (Section 3.1). Then, we dicuss how to adapt the algorithm to Cubical Agda’s type system (Section 3.2). Finally, we show how we can leverage Cubical Agda to produce internal proofs of correctness for the individual components of proof repair (Section 3.3)—something that was not possible in Coq, and was not even done metatheoretically for any examples. We will illustrate this approach on our case studies and discuss the tradeoffs relative to Coq in Sections 4 and 5, respectively.
@@ 3.1. Proof Repair in Coq
Pumpkin Pi leverages this structure to do repair from one type to a mathematically equivalent type. The key insight is that, by Lambek’s theorem, any equivalence between types A and B can be decomposed into components that talk only about A and only about B respectively (Ringer, 2021). Functions and proofs can be unified with applications of these components, making repair a simple proof term transformation (Ringer et al., 2021).
The original Pumpkin Pi work calls each such decomposed equivalence a configuration, comprising pairs of the form ((DepConstr, DepElim) , ($\iota$, $\eta$)) for types on both sides of the equivalence. [^4] For the purposes of this paper, $\eta$, dealing with $\eta$-expansions of constructors applied to eliminators, will always be trivial, and thus we will ignore it. DepConstr and DepElim are, respectively, constructors and eliminators for each type. The constructors must generate the elements of the inductive type, and the eliminator must specify how to consume an element produced by the constructors. These constructors and eliminators take the shape of the original type A, even if the repaired type B has a different shape. To use an example from the Pumpkin Pi paper, for the type of unary naturals in Figure 2, we could give the following dependent constructors:
  Definition depConstrNatZero : Nat.
  Definition depConstrNatSuc : Nat -> Nat.
Then, to do repair to the binary naturals, we need to provide dependent constructors for that type. These dependent constructors must correspond across the equivalence. Thus, we arrive at the following constructors:
  Definition depConstrNZero : N.
  Definition depConstrNSuc : N -> N.
  Definition depElimNat : forall (P : nat -> Type),
    (P depConstrNatZero) ->
    (forall n : nat, P n -> P (depConstrNatSuc n)) ->
    forall n : nat, P n.
  Definition depElimN : forall (P : N -> Type),
    (P depConstrNZero) ->
    (forall n : N, P n -> P (depConstrNSuc n)) ->
    forall n : N, P n.
@@
Figure 3. Transformation for repair across $A\simeq B$ in our fragment with configuration ((DepConstr, DepElim), (Eta, Iota)), from previous work (Ringer, 2021).
3.2. Proof Repair in Cubical Agda

### 15 Mar 2024 v3

# Internal and External Views

@@ Abstract.
Proofs in proof assistants like Coq can be brittle, breaking easily in response to changes in the terms and types those proofs depend on. To address this, recent work introduced an algorithm and tool in Coq to automatically repair broken proofs in response to changes that correspond to type equivalences. However, many changes remained out of the scope of this algorithm and tool—especially changes in underlying behavior. We extend this proof repair algorithm so that it can express certain changes in behavior that were previously out of scope. We focus in particular on equivalences between quotient types—types equipped with a relation that describes what it means for any two elements of that type to be equal. Quotient type equivalences can be used to express interesting changes in representations of mathematical structures, as well as changes in the underlying implementations of data structures—two use cases highlighted by our case studies.
[…]
Proof Repair, Cubical Agda, Coq, Quotient Types
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Proof repair (Ringer, 2021) aims to simplify this process by introducing algorithms and tools that fix formal proofs in response to breaking changes. In this paper, we extend proof repair to handle a new class of breaking changes in datatypes not handled by prior work. Prior work introduced a Coq plugin called Pumpkin Pi for proof repair across changes in datatypes that can be described by type equivalences (Ringer et al., 2021). Pumpkin Pi could handle only a limited class of changes corresponding to non-equivalent types by way of expressing those changes as equivalences between sigma types. The power of this was limited, however, and could not represent fundamental changes in behavior in a practical way.
In this work, we extend proof repair to support equivalences between quotient types (Section 2). Recent work in Cubical Agda showed that certain relations describing changes in behavior can be adjusted to equivalences between quotient types (Angiuli et al., 2021). However, Coq lacks quotient types entirely, so one cannot use the original Pumpkin Pi transformation (Section 3) as-is to support this class of changes. Accordingly, we explore two different views of proof repair across quotient type equivalences:
@@
    an internal view using native quotient types in Cubical Agda (Section 4), and
@@
    an external view using setoids in Coq (Section 5).
Our exploration includes two case studies in both Coq in Cubical Agda, neither of which can be handled by prior work:
@@
    We examine two non-isomorphic representations of the integers: (a) the representation used in the Cubical Agda standard library, and (b) the Grothendieck group completion of the natural numbers. We repair the addition function and proofs about it (Section 6.1).
@@
    We implement two variations of queues: (a) those backed by lists and (b) those backed by pairs of lists. These types are not equivalent, but we can construct a quotient type equivalence that describes the change and repair functions and proofs across it (Section 6.2).
Our findings are summarized in Table 1. Our code is available in supplementary material.  [^1] https://github.com/InnovativeInventor/proof-repair-quotients
Table 1. Comparison of proof repair across quotient type equivalences in Cubical Agda and Coq.
Feature Cubical Agda Coq
Quotient Types Internal External via Setoids
Repair Transformation Straightforward Extended
Correctness Some Internal External
Automation None Prototype
An Internal View in Cubical Agda
For our internal view, we adapt the transformation at the heart of the algorithm for proof repair across type equivalences to a fragment of cubical type theory, the type theory at the heart of Cubical Agda. The tradeoffs, in summary:
@@
    Quotient Types: Internal. Cubical Agda supports quotient types natively, by way of a particular higher-inductive type. Thus, we do not need to do anything special to represent quotient type equivalences; they are just equivalences between quotient types.
@@
    Repair Transformation: Straightforward. A consequence of native support for quotients is that we can adapt the transformation at the heart of proof repair across type equivalences from prior work, and use it in a straightforward way for proof repair across quotient type equivalences over our supported fragment of cubical type theory without any extensions.
@@
    Correctness: Some Internal. We can state and prove the correctness of repaired proofs internally in Cubical Agda, using cubical’s dependent path equality. Thanks to this, we are able to prove the correctness of repaired induction principles for the first time, something that was not possible in Pumpkin Pi using Coq. Still, actually composing these correctness proofs proves hard thanks to the proof-relevant notion of equality in cubical.
  4. (4)
    Automation: None. Automation in Cubical Agda is difficult due to engineering limitations and type-theoretic challenges. Our proof repair case studies in Cubical Agda are manual.
An External View in Coq
For our external view, we represent quotient types using Coq’s setoid machinery, and extend the same proof transformation to support the newly generated equality proof obligations. We also extend Pumpkin Pi with new prototype automation. The tradeoffs, in summary:
  1. (1)
    Quotient Types: External via Setoids. Coq lacks internal support for quotient types. We represent quotient types externally using Coq’s setoid machinery, and we represent quotient type equivalences as setoid equivalences.
  2. (2)
    Repair Transformation: Extended. One consequence of the lack of support for quotient types is that we must extend the proof repair algorithm to support the setoid machinery. This involves adding new cases corresponding to equality proofs and types.
  3. (3)
    Correctness: External. As in prior work, since Coq lacks univalence, we can only specify correctness by way of previously stated correctness theorems on paper inside of a univalent metatheory. We cannot state and prove them in general in Coq without introducing axioms.
  4. (4)
    Automation: Prototype. Thanks to Coq’s rich plugin infrastructure, we implement a prototype extension to the Pumpkin Pi proof repair automation in just $763$ lines of code to handle the new equality cases of the extended repair algorithm.
@@ 2. Problem Definition
We expand this problem to allow for the set of pairs of equivalent types $A$ and $B$ that we can repair proofs over to include an extremely useful class of changes in types not describable in Coq—those that can be described by quotient type equivalences (Section 2.1). For simplicity, we start by considering this problem in a fragment of cubical type theory which supports quotient types directly. In such a context, we can internally and formally state what it means for any given repaired proof defined over $B$ to correctly correspond to the original proof over $A$—in terms of dependent path equality (Section 2.2). To implement this, we adapt the proof term transformation implemented in Coq for the Pumpkin Pi plugin (Section 3). We dually adapt this to support a fragment of Cubical Agda that has native support for quotient types (Section 4) and to support an extended fragment of Coq that represents quotient types using Coq’s setoid machinery (Section 5).
@@ 2.1. Scope: Quotient Type Equivalences
In this paper, we perform proof repair across quotient type equivalences. We will describe what makes these quotient type equivalences soon. But in general, a type equivalence between two types $A$ and $B$ is an isomorphism (a pair of functions that are mutual inverses) satisfying a particular coherence property (Univalent Foundations Program, 2013). It is possible to form an equivalence from any isomorphism.
@@
Like Pumpkin Pi in Coq, we consider changes in datatypes that can be described by these type equivalences. The key difference between our work and Pumpkin Pi is that we consider a broader class of changes than type equivalences in Coq can encode. The trick is to encode changes in datatypes as equivalences between quotient types—types equipped with an equivalence relation describing what makes two elements of that type “the same.”
Quotient types are supported internally in cubical type theory by way of higher inductive types. We denote quotient types as $A/R$ for a type $A$ and a relation $R$. An element of this quotient is the equivalence class of any $a$ of type $A$. Given two elements $[a_{1}]$ and $[a_{2}]$, we have that $[a_{1}]\equiv[a_{2}]$ whenever $R\ a_{1}\ a_{2}$—even when it is not true that $a_{1}\equiv a_{2}$.
@@
To give a familiar example of a quotient type, we can construct the natural numbers mod 2, $\mathbb{N}/2$. Two naturals are in the same equivalence class if they have the same parity. The resulting type has two equivalence classes, and so every element of the type is equal to either [ 0 ] or [ 1 ]. For another element, like [ 2 ], we have a proof of equality between [ 0 ] and [ 2 ], since both are even. This resulting type is then isomorphic to the booleans via the map [ 0 ] $\mapsto$ false, [ 1 ] $\mapsto$ true, and this can be used to construct a type equivalence.
@@
In this paper, we further restrict ourselves to equivalences between set quotients, which have the additional constraint that all proofs of equality between elements of the quotient must themselves be equal to each other. In homotopy type theory parlance, $A/R$ must be an h-set; in broader dependent type theory parlance, uniqueness of identity proofs must provably hold for $A/R$ [^2] While homotopy type theory makes it possible to define types for which identity proofs are not unique, reasoning about equalities between those types can be challenging, so we defer reasoning about them to later work.
Quotient type equivalences let us naturally express changes in underlying implementation that break equivalence. They are useful for mathematics as well as for changes in implementation of datatypes. For example, in Section 6.2 we will show how to use quotient type equivalences to repair functions and proofs between two non-isomorphic implementations of a queue: one backed by a single list, and one backed by a pair of lists. For the latter representation, we will implement an efficient dequeueing function that operates by removing an element from the front of the second list of the pair, and an efficient enqueueing function that works by appending to the front of the first list in the pair. We will then repair proofs between these inefficient and efficient representations.
Extending the Pumpkin Pi transformation to support quotient type equivalences is what makes repair across this change possible for the first time. In Pumpkin Pi, some changes more interesting than equivalences could be expressed as equivalences between sigma types; we could try that approach for quotients, but it would come with severe disadvantages. For example, we could represent queues backed by pairs of lists as the sigma type:
$\sum_{x\>:\>\text{List }A\>\times\>\text{List }A}isCanonical\ x$
where $isCanonical\ x$ is a proposition stating that $x$ is the canonical representative of its equivalence class. This type is equivalent to our type of one list queues. However, this type has only one representative of each equivalence class present within it. With only one representative of each equivalence class present in the type, the efficient implementations of enqueue and dequeue would not be possible. Quotient types empower us to repair functions and proofs to operate over the efficient representation.
@@ 2.2. Goal: Dependent Path Equality
In cubical type theory, we can state what it means for a repaired proof to be correctly related to the original proof. This is because cubical type theory is univalent (Cohen et al., 2018; Angiuli et al., 2017)—equivalence is equivalent to equality (Univalent Foundations Program, 2013). [^3] In cubical type theory, univalence is not an axiom, but rather follows computationally from more primitive constructs. We take advantage not of univalence directly, but of these more primitive constructive constructs.
We define correctness in terms of a generalization of path equality, the primitive notion of propositional equality in cubical type theory. If two terms a and b of the same type A are path equal to each other, we can produce a term of type a $\equiv$ b.
Our two-element types from before provide an example of path equality. Using univalence, the isomorphism between $\mathbb{N}/2$ and the booleans can be turned into an equality $\mathbb{N}/2$equiv bool.
Path equality is actually an instantiation of the more powerful notion of dependent path equality, for equality between terms with different types (that is, a kind of heterogeneous equality). In this paper, we use the Cubical Agda syntax to denote dependent path equality. That is, two terms a of type A and b of type B are dependently path equal if we can produce a term of type:
@@
where p is a path between A and B. Non-dependent path equality $\equiv$ is defined in terms of dependent path equality at the identity path. For example, let p be our proof that $\mathbb{N}/2$equiv bool. Then,
@@
We can state correctness of a repaired proof relative to the corresponding original version by stating the correct PathP type—the one between the original term and the repaired term at the path between their types. The type PathP p [ 0 ] false we described above gives such an example: if we repair [ 0 ] to false across $\mathbb{N}/2\equiv$ bool, then this type states that this repair was correct.
We will describe how to define correctness more generally using PathP when we detail our internal approach in Cubical Agda in Section 4. Our external approach in Coq in Section 5 will rely solely on metatheoretical statements of correctness from Pumpkin Pi which, consistently with prior work, are not proven. One other thing to note that is also consistent with prior work is that correctness of a repaired proof relies not only on the right PathP type being inhabited, but also on the new version of the proof no longer containing references to the old version of the proof. This we do not state internally in either Cubical Agda or Coq, but it should still hold metatheoretically for repair to be correct. This is also the primary way in which proof repair in cubical type theory diverges from (dependent) transport: using transport to move proofs across equalities forces both equivalent types to eternally remain in the codebase. Repair lets us remove the old type.
3. Approach: Proof Term Transformation
To extend proof repair to support quotient type equivalences, we adapt the transformation corresponding to Pumpkin Pi’s proof repair algorithm in two different ways:
  1. (1)
    to support a fragment of cubical type theory for Cubical Agda (Section 4), and
  2. (2)
    to support an external notion of quotient types in Coq (Section 5).
The former corresponds to our internal view, while the latter corresponds to our external view.
@@ 3. Approach: Proof Term Transformation
[…]
@@
Pumpkin Pi calls each such decomposed equivalence a configuration, comprising pairs of the form ((DepConstr, DepElim) , ($\iota$, $\eta$)) for types on both sides of the equivalence. [^4] For the purposes of this paper, $\eta$, dealing with $\eta$-expansions of constructors applied to eliminators, will always be trivial, and thus we will ignore it. DepConstr and DepElim are, respectively, constructors and eliminators for each type. The constructors must generate the elements of the inductive type, and the eliminator must specify how to consume an element produced by the constructors. These constructors and eliminators take the shape of the original type A, even if the repaired type B has a different shape. To use an example from the Pumpkin Pi paper, for the type of unary naturals in Figure 2, we could give the following dependent constructors:
  Definition depConstrNatZero : Nat.
  Definition depConstrNatSuc : Nat -> Nat.
To repair to the binary naturals, we need to provide dependent constructors N. These dependent constructors must correspond across the equivalence. Thus, we arrive at the following constructors:
  Definition depConstrNZero : N.
  Definition depConstrNSuc : N -> N.
  Definition depElimNat : forall (P : nat -> Type),
    (P depConstrNatZero) -> (forall n : nat, P n -> P (depConstrNatSuc n)) ->
    forall n : nat, P n.
  Definition depElimN : forall (P : N -> Type),
    (P depConstrNZero) -> (forall n : N, P n -> P (depConstrNSuc n)) ->
    forall n : N, P n.
@@
Figure 3. Transformation for repair across $A\simeq B$ with configuration ((DepConstr, DepElim), (Eta, Iota)), from previous work (Ringer, 2021). Our work adapts and extends this transformation.
We adapt this transformation for our work on proof repair across quotient type equivalences in both Cubical Agda and Coq. Note that Pumpkin Pi implements significant automation to run this transformation on existing proof terms in Coq. It does not support quotient type equivalences at all, however, since Coq does not support quotient types. Our work in Cubical Agda is manual, but supports quotient types internally; our work in Coq includes prototype automation that extends Pumpkin Pi’s automation directly, but relies on an external representation of quotient types in Coq.
4. Internal View: Proof Repair with Quotient Types in Cubical Agda
Cubical Agda supports quotient types internally. This means that, so long as we can adapt the proof term transformation to a fragment of Cubical Agda’s underlying type theory, we get proof repair across quotient type equivalences “for free,” since quotient type equivalences are just regular type equivalences (Section 4.1). Luckily, we find that adapting the algorithm to a fragment of Cubical Agda’s type system is straightforward (Section 4.2). In addition, we can leverage Cubical Agda’s dependent path equality to produce internal proofs of correctness for the individual components of proof repair (Section 4.3)—something that was not possible in Coq, and was not even done metatheoretically for any examples. Of course, one thing is notably missing from our approach in Cubical Agda: automation (Section  4.4).
4.1. Quotient Types: Internal
In Cubical Agda, set quotient types are encoded internally as higher inductive types:
    [_] : (a : A) normal-$\rightarrow$ A / R